OpenStudy (anonymous):

PLEASE HELP WILL MEDAL AND FAN AND EVERYTHING IM DESPRATE Which is not a solution to the equation -3x + y =5 ? A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8)

OpenStudy (anonymous):

change it to y=3x+5 and put in (-4,-7) -7=3(-4)+5 =-12+5 =-7 correct now do the same with the other points to find the odd one which is not equal both sides

OpenStudy (anonymous):

What? @jamaica25

OpenStudy (anonymous):

I will medal and fan please helppp

OpenStudy (anonymous):

i would say C

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@Rocky_

OpenStudy (anonymous):

np